OpenStudy (anonymous):

what is 2 4/3 equal to? the answer must be in square root form

OpenStudy (jdoe0001):

hmmm sounds what do they mean by "square root form"? that could mean anything

OpenStudy (anonymous):

like the answer is an exponent and then there's a radical and then it has a base in the radical

OpenStudy (anonymous):

\[2^\frac{ 4 }{3 }\]?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

actually, got the colors backwards there =) \(\Large { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ 2^{\frac{{\color{blue}{ 4}}}{{\color{red}{ 3}}}}\implies ? }\)

OpenStudy (anonymous):

so ^4 sqrt 2^3 ?

OpenStudy (anonymous):

and then what's after that

OpenStudy (jdoe0001):

hmmm the numerator is the exponent the denominator is the root

OpenStudy (jdoe0001):

\(\large { \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ 2^{\frac{{\color{blue}{ 4}}}{{\color{red}{ 3}}}}\implies \sqrt[{\color{red}{ 3}}]{2^{{\color{blue}{ 4}}}}\implies \sqrt[3]{2\cdot 2\cdot 2\cdot 2\cdot 2}\implies \sqrt[3]{2^3\cdot 2^1}\implies 2\sqrt[3]{2} }\)

OpenStudy (jdoe0001):

hmm actuallly, I have an extra 2 there =) \(\large { \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ 2^{\frac{{\color{blue}{ 4}}}{{\color{red}{ 3}}}}\implies \sqrt[{\color{red}{ 3}}]{2^{{\color{blue}{ 4}}}}\implies \sqrt[3]{2\cdot 2\cdot 2\cdot 2}\implies \sqrt[3]{2^3\cdot 2^1}\implies 2\sqrt[3]{2} }\)

OpenStudy (anonymous):

oh so it'd be 3 sqrt 16